# LLM-based feature generation from text for interpretable machine learning

### v1

@@ Abstract
Existing text representations such as embeddings and bag-of-words are not suitable for rule learning due to their high dimensionality and absent or questionable feature-level interpretability. This article explores whether large language models (LLMs) could address this by extracting a small number of interpretable features from text. We demonstrate this process on two datasets (CORD-19 and M17+) containing several thousand scientific articles from multiple disciplines and a target being a proxy for research impact. An evaluation based on testing for the statistically significant correlation with research impact has shown that LLama 2-generated features are semantically meaningful. We consequently used these generated features in text classification to predict the binary target variable representing the citation rate for the CORD-19 dataset and the ordinal 5-class target representing an expert-awarded grade in the M17+ dataset. Machine-learning models trained on the LLM-generated features provided similar predictive performance to the state-of-the-art embedding model SciBERT for scientific text. The LLM used only 62 features compared to 768 features in SciBERT embeddings, and these features were directly interpretable, corresponding to notions such as article methodological rigor, novelty, or grammatical correctness. As the final step, we extract a small number of well-interpretable action rules. Consistently competitive results obtained with the same LLM feature set across both thematically diverse datasets show that this approach generalizes across domains.
@@ 1 Introduction
Large language models (LLMs) have seen a significant increase in use in academic and commercial spheres alike, and the wide availability of open-weight models such as LLama2 [18] makes LLMs suitable for use as a component in machine learning workflows. These models are commonly used for text summarization, classification, or natural language generation. In this work, we evaluate LLMs on the task of feature generation from text and then show how these newly generated features can be used for rule learning.
As the specific use case, we apply LLM-based feature generation on the task of scholarly document quality prediction. First, we use an LLM to assess a large number of scientific article abstracts, evaluating them according to multiple criteria such as methodological rigor, grammatical correctness, article novelty, or text accessibility.
Example. Our LLM feature generator applied to a hypothetical initial version of this articles’ abstract generated features such as rigor=high (on a three point scale low/medium/high), grammar=0 (on binary scale, where 0 is free of errors and 1 is contains errors), novelty=low, accessibility=high, replicability=0 (not good). Action rules were learnt on a training dataset with similarly LLM-annotated features, and a target corresponding to human expert assessment of article quality. An applicable action rule found for our abstract is: $r_{15}:\text{novelty}=(\text{low}\rightarrow\text{high})\wedge\text{replicability}=(0\rightarrow 1)\Rightarrow\text{evaluation}=(\text{avg}\rightarrow\text{best}).$ This rule, which used only two of the available features, indicates that in its current form, the article is predicted to receive an average expert evaluation, but if applicability and novelty are improved, the probability of the evaluation increasing to a higher rating will rise. [^1] The full listing of $r_{15}$ in Section 3.5) also includes uplift, an indicator of probabilistic nature of this recommendation. Consequently, we could modify the abstract to emphasize that LLM-based features have not yet been used in rule learning and that our use of an open LLM fosters replicability compared to prior work that used closed commercial LLMs.
The first contribution of our work is in a multi-pronged evaluation of the quality of LLM-generated features, including their agreement with domain expertise, analysis of interpretability, and predictive performance. As reference methods, we use both state-of-the-art but ”black-box” embedding-based methods as well as the partly interpretable BoW representation. We also demonstrate the effect of the fusion of LLM-based and BoW features.
@@ 2 Methods
In this research we employ the publicly available open-source LLama2 model and use it to extract several features from scientific article abstracts. These features are then used as inputs for several machine-learning algorithms, and their predictive performance is measured against alternative forms of text representations, such as word embeddings or TF-IDF matrix.
@@ 2.1 Input data
The data used for experiments in this work are two datasets that contain scientific article titles, abstracts, and a target variable.
The CORD-19 dataset is the first dataset of interest. It is primarily composed of scientific articles related to coronaviruses that span from the 1970’s to 2020, the beginning of the COVID-19 outbreak. The target variable in this dataset represents either a low or high citation rate. This variable was computed in previous research by [4] by dividing the number of citations obtained from OpenCitations and Web of Science by the age of the publication in years and then assigning 0 to articles below the median and 1 to articles above the median. In order to reduce the computational complexity, the dataset size was reduced to 3000 articles and was balanced with respect to the target variable.
@@
The M17+ dataset [5] is also used in this work’s experiments. The dataset consists of 7710 articles from Czech research institutions that were graded by a jury of experts in a program subsidized by the government of the Czech Republic, similar to the U.K. Research Excellence Framework 2021 (REF) that has been the subject of previous research in regards to grade prediction of articles by machine-learning methods [17]. The target variable in this dataset is an ordinal variable ranging from 1 to 5, with 1 meaning that the article is of world-class quality and impact, while 5 means that the article is of mediocre quality and impact. The dataset was downsampled from 7710 articles to 2000 with respect to the target variable. This resulted in an even distribution of 400 articles belonging to each class.
@@
Both datasets contained several columns beside the article title, abstract, and target variable, such as journal name, impact factors (IF), or article influence score (AIS). The final version of both datasets contained only 3 columns - title, abstract, and target.
2.2 Feature derivation
Based on the article abstracts, several features were extracted using the non-finetuned LLama2 13B GPTQ Chat model hosted on our own hardware. The features generated in the experiment were based on prior research from various authors that mainly focused on inspecting the relationship between citation rate/count and respective features that were derived by other means than LLM inference.
The features are defined as outlined in Table 1.
@@ 2.2 Feature derivation
Reproducibility Assessed the authors’ mention of the reproducibility of results presented in the abstract {no, yes}
@@
Table 1: Generated features for article abstracts, a total of 62 features were added to the datasets.
@@
The research type and discipline features were divided into multiple binary variables indicating the articles belonging to respective research types or disciplines. Each article was allowed to belong into more than one discipline and research.
@@
We used the LLama2 13B GPTQ Chat [^2] https://huggingface.co/TheBloke/Llama-2-13B-chat-GPTQ model through the transformers library API [19], which allowed us to set model hyperparameters so that the models produced consistent results.
2.3 LLM feature quality validation
In this research, we generate various categorical features using LLMs. Evaluation of LLM output is difficult, especially when dealing with text characteristics that can be subjective, i.e., methodological rigor or accessibility. Currently, the most reliable way of validating LLM output is human evaluation, which is also the most expensive and can suffer from having high variance and instability, depending on the human evaluators [6].
We evaluate the generated features in two ways. In the standard quantitative approach, we observe the effect of the feature on predictive performance. In the qualitative approach, we suppose there to be certain relationships between the generated features and the target variables. For example, we suppose that articles with higher methodological rigor will be, on average, graded higher than articles with lower rigor. The qualitative validation is based on checking whether we observe the corresponding correlations between the generated feature and the target.
2.4 Alternative text data representation
In order to compare the performance of machine-learning algorithms trained on LLM-generated features, we also employed two of the common techniques of text representation to compare the predictive performance.
Bag-of-Words (BoW) is a method for representing text as an unordered collection of terms. In this context, we use the term frequency-inversed document frequency (TF-IDF) matrix to obtain term weights. The TF-IDF matrix was constructed for 1500 terms, including unigrams, bigrams, and trigrams. To filter out domain-specific stopwords and irrelevant terms, the minimum document frequency was set to 0.005, and the maximum document frequency was set to 0.80 for the TF-IDF terms.
Word Embeddings is employed through the use of state-of-the-art language models for scientific text SciBERT [3]. SciBERT was trained on a large amount of scientific data and significantly outperformed the base BERT model [7]. The embeddings produced by this model are vectors of length 768.
2.5 Machine-learning algorithms
Scikit-learn python library was used as a source for a range of models [11]. Specifically, we tested the predictive performance of the Random Forest Classifier, Gradient Boost Classifier, Ada Boost Classifier, Support Vector Machines, and Logistic Regression. For the ensemble models, the best models were chosen from the hyperparameter space in Table 2 using grid search. Out of the aforementioned models, the Gradient Boost Classier consistently showed the best results. Therefore, we used only this model for subsequent experiments with LLM and BoW features.
Hyperparameter Values
Max Depth 10, 20, 30, 40, 45, 50
Max Features log2, sqrt, 30
Min Samples Leaf 30, 40, 50, 60, 70, 90
Min Samples Split 20, 30, 50, 60, 90
Number of Estimators 100, 200, 400, 1000
Table 2: Hyperparameter space for ensemble models
AutoGluon is a deep-learning API used for training neural networks on a combination of tabular and image, text, or audio data [16]. In this study, we employed the AutoGluon Tabular model with title, abstract and LLM-generated features as input. We tested both regressor and classifier configurations. The classifier configuration provided better results in all observed metrics and was thus used as the only variation of the AutoGluon model.
@@ 2.7 Feature subsets
For the comparisons of predictive performance in a classification task, various feature subsets were generated for both the CORD-19 and M17+ datasets (Table 3). For the sake of reproducibility and comparability, both CORD-19 and M17+ sets were split to train and test sets in an 80:20 ratio with fixed random seed before the application of feature subsetting. This resulted in the same articles being tested each time, only with different features:
text only dataset uses only title and abstract columns from original datasets as input into the AutoGluon Tabular model.
BoW only is a feature set where the only features used are bag-of-words terms weighted by their TF-IDF score along with the target variable. Due to computational constraints, 1500 terms were considered for unigrams, bigrams, and trigrams. These sets have 1501 columns, including the target variable.
SciBERT embeddings only is a feature set where, besides target variables, each abstract was embedded using the SciBERT model. The dataset has 796 columns, including the target variable.
LLM-features only dataset represents CORD-19 and M17+ sets, where every column beside the target variable is LLM-generated. These datasets have 63 columns – 62 feature columns and 1 target variable.
BoW + LLM-generated features is a dataset of 1500 TF-IDF-weighted terms and all 62 feature columns from LLM inference, totaling 1564 columns with the target variables, including the target variable.
@@
Table 3: Overview of feature combinations for experiments on the classification tasks, CORD-19 and M17+. The number of features does not include the target variable. $\dagger$ The two features were text data in abstract and title, vectorization was performed internally in AutoGluon and is not separately reported. *Consistently showed the best results and was used for the final experiments.
Feature Combination No. of Features Used Models
text only $2$ AutoGluon Tabular
BoW only $1500$ SVM, LogReg, GradientBoost*, AdaBoost, RandomForest
SciBERT embeddings only $768$ LSTM Neural Network
LLM features only $62$ SVM, LogReg, GradientBoost*, AdaBoost, RandomForest
BoW + LLM-generated features $1562$ SVM, LogReg, GradientBoost*, AdaBoost, RandomForest
text + LLM-generated features $64$ AutoGluon Tabular
@@ 2.8 Action rules
An action rule $r_{a}$ can be decomposed into two classification rules, $r_{undesired}$ and $r_{desired}$: one that predicts the undesired state of the target (before intervention), and another that predicts the desired state (after intervention), represented as $r_{undesired}\rightarrow r_{desired}$. Uplift is a measure used to predict the incremental response to an action [12], see Equation 7.
@@
During the generation of action rules, it is necessary to address the issue of generating a large number of highly similar rules. The approach dominant action rule to reducing action rules is inspired by closed association rules [10]. This concept is analogous to dominant action rules in that closed rules cannot be extended without a loss of support, which can be likened to the reduction in uplift seen in dominant rules. A dominant action rule is one that cannot be further expanded with additional conditions to increase its uplift. For example, the rule $r_{4}$ (see Equation 8) is dominant over the longer rule $r_{5}$. In this specific case, the conditions in rule $r_{4}$ are not extended in $r_{5}$ in a way that increases uplift. Instead, the addition of another condition in $r_{5}$ results in a decreased uplift, demonstrating the dominance of $r_{4}$ over $r_{5}$.

### v2

@@ Abstract
Traditional text representations like embeddings and bag-of-words hinder rule learning and other interpretable machine learning methods due to high dimensionality and poor comprehensibility. This article investigates using Large Language Models (LLMs) to extract a small number of interpretable text features. We propose two workflows: one fully automated by the LLM (feature proposal and value calculation), and another where users define features and the LLM calculates values. This LLM-based feature extraction enables interpretable rule learning, overcoming issues like spurious interpretability seen with bag-of-words. We evaluated the proposed methods on five diverse datasets (including scientometrics, banking, hate speech, and food hazard). LLM-generated features yielded predictive performance similar to the SciBERT embedding model but used far fewer, interpretable features. Most generated features were considered relevant for the corresponding prediction tasks by human users. We illustrate practical utility on a case study focused on mining recommendation action rules for the improvement of research article quality and citation impact.
@@ 1 Introduction
Large language models (LLMs) have seen a significant increase in use in academic and commercial spheres alike, and the wide availability of open-weight models such as Llama2 [32] makes LLMs suitable for use as a component in machine learning workflows. These models are commonly used for text summarization, classification, or natural language generation.
This article explores whether large language models (LLMs) could address this by extracting a small number of interpretable features from text. We propose two workflows, one fully automated, where LLM both proposes suitable features and computes their values, and a second one where the feature names are user-specified and LLM only generates their values. We further show that LLM-based feature extraction can be used as an input for rule learning and has well-interpretable results. We evaluate this process on two datasets (CORD-19 and M17+) from the scientometric domain and three additional datasets from other domains: customer intent from the banking domain, hate speech and food hazard.
To illustrate the process, consider the scholarly document quality prediction use case. First, we use an LLM to assess a large number of scientific article abstracts, evaluating them according to multiple criteria. These criteria can either be LLM-generated or user-specified, as in the example below.
Example. Our LLM feature generator applied to a hypothetical initial version of this articles’ abstract generated features such as rigor=high (on a three point scale low/medium/high), grammar=0 (on binary scale, where 0 is free of errors and 1 is contains errors), novelty=low, accessibility=high, replicability=0 (not good). Action rules were learnt on a training dataset with similarly LLM-annotated features, and a target corresponding to human expert assessment of article quality. An applicable action rule found for our abstract is: $r_{15}:\text{novelty}=(\text{low}\rightarrow\text{high})\wedge\text{replicability}=(0\rightarrow 1)\Rightarrow\text{evaluation}=(\text{avg}\rightarrow\text{best}).$ This rule, which used only two of the available features, indicates that in its current form, the article is predicted to receive an average expert evaluation, but if replicability and novelty are improved, the probability of the evaluation increasing to a higher rating will rise. [^1] The full listing of $r_{15}$ in Section 3.4) also includes uplift, an indicator of probabilistic nature of this recommendation. Consequently, we could modify the abstract to emphasize that LLM-based features have not yet been used in rule learning and that our use of an open LLM fosters replicability compared to prior work that used closed commercial LLMs.
The first contribution of our work is in a multi-pronged evaluation of the quality of LLM-generated features, including their agreement with domain expertise, analysis of interpretability, and predictive performance. As reference methods, we use both state-of-the-art but “black-box” embedding-based methods as well as the partly interpretable BoW representation. We also demonstrate the effect of the fusion of LLM-based and BoW features.
@@ 2 Methods
In this research, we employ the publicly available open-source Llama2 model and use it to extract several features from scientific article abstracts. These features are then used as inputs for several machine-learning algorithms, and their predictive performance is measured against alternative forms of text representations, such as word embeddings or TF-IDF matrix.
@@ 2.1 Input data
Five datasets were used for experiments in this work. Two datasets (CORD-19 and M17+) contain scientific article titles, abstracts, and a target variable related to research impact. To demonstrate the universality of the research, three additional datasets from fields other than scientometrics are utilized: BANKING77222https://paperswithcode.com/dataset/banking77, Hate Speech [^3] https://paperswithcode.com/dataset/hate-speech and Food Hazard [^4] https://food-hazard-detection-semeval-2025.github.io/.
Scientometric datasets
[…]
From both datasets, we use article abstracts as the main source of information. The use of abstracts improves the applicability of our method, as abstracts are more easily available.
@@
The M17+ dataset [6] is also used in this work’s experiments. The dataset consists of 7,710 articles from Czech research institutions that were graded by a jury of experts in a program subsidized by the government of the Czech Republic, similar to the U.K. Research Excellence Framework 2021 (REF) that has been the subject of previous research with regards to grade prediction of articles by machine-learning methods [31]. The target variable in this dataset is an ordinal variable ranging from 1 to 5, with 1 meaning that the article is of world-class quality and impact, while 5 means that the article is of mediocre quality and impact. The dataset was downsampled from 7,710 articles to 2,000 with respect to the target variable. This resulted in an even distribution of 400 articles belonging to each class.
@@
Both scientometric datasets contained several columns beside the article title, abstract, and target variable, such as journal name, impact factors (IF), or article influence score (AIS). The final version of both datasets contained only 3 columns - title, abstract, and target.
Other datasets
The BANKING77 dataset [^5] https://paperswithcode.com/dataset/banking77 provides a fine-grained set of intents in the banking domain. It comprises 13,083 customer service queries labeled with 77 intents. This dataset focuses on fine-grained single-domain intent detection and is commonly used for intent classification tasks [7].
The Hate Speech dataset [^6] https://paperswithcode.com/dataset/hate-speech consists of sentences extracted from Stormfront, a white supremacist forum. A total of 10,568 sentences have been manually annotated as conveying hate speech or not. This dataset is utilized for research in hate speech detection and related tasks [9].
The Food Hazard dataset [^7] https://food-hazard-detection-semeval-2025.github.io/ is a part of the Semeval-2025 challenge and contains 6,644 food incident reports containing information about product recalls from regulatory bodies along with 4 target variables that correspond to 4 classification tasks. We chose the task of predicting the ’hazard-category’ target variable.
All datasets were preprocessed to retain only the relevant textual data and target variables necessary for our experiments.
2.2 Machine-learning algorithms and their hyperparameters
The Scikit-learn Python library was used as a source for a range of models [24]. Specifically, we tested the predictive performance of the Random Forest Classifier, Gradient Boost Classifier, Ada Boost Classifier, Support Vector Machines, and Logistic Regression on CORD-19 and M17+ datasets. For the ensemble models, the best models were chosen from the hyperparameter space in Table 1 using grid search. Out of the aforementioned models, the Gradient Boost Classifier consistently showed the best results. Therefore, we used only this model for subsequent experiments with LLM and BoW features on M17+ and CORD-19 datasets. For experiments on BANKING-77, Hate Speech, and Food Hazard datasets, which were several times larger than M17+ and CORD-19, we used the Random Forest Classifier, which resulted in a very similar performance to the Gradient Boost Classifier but was faster.
Settings for TF-IDF, which served as one of the baselines, are in the provided GitHub repository.
Hyperparameter GradientBoost Values RandomForest Values
Max Depth 10, 20, 30, 40, 45, 50 3, 5, 10, None
Max Features log2, sqrt, 30
Min Samples Leaf 30, 40, 50, 60, 70, 90 1, 2, 5
Min Samples Split 20, 30, 50, 60, 90 2, 5, 10
Number of Estimators 100, 200, 400, 1000 50, 100, 200
Table 1: Hyperparameter space for ensemble models
In addition to scikit-learn, for experiments on the two scientometric datasets we used the AutoGluon library. AutoGluon is a deep-learning API used for training neural networks on a combination of tabular and image, text, or audio data [30]. In this study, we employed the AutoGluon Tabular model with title, abstract and LLM-generated features as input. We tested both regressor and classifier configurations. The classifier configuration provided better results in all observed metrics and was thus used as the only variation of the AutoGluon model.
2.3 LLM-based feature generation based on user-set feature list
The first method of LLM-assisted feature generation is based on one-shot generation with user-constructed prompts for manually selected features. This method was used for the extraction of features for CORD-19 and M17+ datasets only. The features shown in Table 2 were manually selected in this experiment by the authors. The selection of feature names was based on our knowledge of the subject domain from previous publications [5, 11] that also focused on inspecting the relationship between article content and research impact.
@@ 2.3 LLM-based feature generation based on user-set feature list
Replicability Assessed the authors’ mention of the reproducibility of results presented in the abstract {no, yes}
@@
Table 2: User-specified features for scientometric datasets. These features were used for LLM-based generation of a total of 62 features (3 ordinal, 59 binary) from article abstracts in M17+ and CORD-19 datasets.
@@
The user-defined features included several ordinal features, such as rigor. The research type and discipline features were originally multinomial and were divided into multiple binary variables indicating the articles belonging to respective research types or disciplines. Each article was allowed to belong to more than one discipline and research.
Listing 1 shows an example of a user-constructed prompt used for LLM-based generation of the value of a specific feature (here ’rigor’) for a given data instance (a research paper abstract).
Listing 1: Example prompt for LLM-based determination of the value of a specific feature
You are a categorization assistant. Your job will be to assign a certain characteristic to a research paper based on its abstract.
In this instance you will assess the methodological rigor of the research.
Methodological rigor represents the soundness of logic presented in the article abstract by its authors.
You will choose between three levels of rigor: low, medium and high. Low being the least rigorous and high being the most.
Be concise, no explanation is to be provided. Your answer will consist of an answer in plain json format and nothing else like so:
""rigor"": ""value""
Abstract to be evaluated:
<abstract>
@@
We used the Llama2 13B GPTQ Chat [^8] https://huggingface.co/TheBloke/Llama-2-13B-chat-GPTQ model through the transformers library API [33], which allowed us to set model hyperparameters so that the models produced consistent results. The hyperparameter do_sample in the model.generate() method was set to False, as that ensured deterministic generation by always selecting the most likely next token, avoiding randomness in the output.
2.4 LLM-based feature generation with automatic feature discovery
In contrast to the previous method, which used LLMs solely for the generation of feature values, this section introduces an enhanced automated workflow leveraging LLMs for feature discovery – determining the names of features to extract. This method further reduces the need for human intervention and was applied to all five datasets.
The core of the approach is a new prompt that contains essential details about the dataset, including its name, a concise description, the text column name, the target column name, and a definition of the target variable. The full prompt is depicted in Figure 4 (Appendix A). As a variable part of the prompt, the model receives a representative sample of rows from the dataset. Based on this information, the model autonomously discovers relevant features for extraction along with corresponding prompts designed explicitly to facilitate their extraction.
The subsequent phase for generating the value of a specific feature for a given data instance is done again automatically with an LLM prompt. The example of a feature generation prompt is depicted in Figure 5 (Appendix A). The difference from the approach described in the previous Section 2.3 and exemplified in Listing 1 is that now not only is the feature value generated (determined, computed) automatically but also the prompt used to generate the feature value is created automatically for a given feature type.
For comparison with the method introduced in the previous subsection, examples of both levels of the automation of feature generation are included in Tables 5-7 in Section 3.2 covering LLM-augmented datasets. Additionally, a direct comparison in terms of predictive performance (SHAP values) is included in the Results section in Figure 1.
In our experiments, automated feature discovery was performed using the model gpt-4o-2024-08-06 (temperature: 0, top_p: 0.9). GPT-4o, introduced by OpenAI in May 2024, is a multimodal large language model capable of processing and generating text, images, and audio. We used 40 randomly selected rows as part of the customization of the feature selection prompt. Following the discovery phase, feature generation was executed using the OpenAI Batch API with the model gpt-4o-mini-2024-07-18 (temperature: 0, top_p: 0.9). To optimize cost efficiency, a 24-hour processing window was chosen.
@@ 2.6 Feature subsets
To compare predictive performance in the classification task, we generated various feature subsets.
For the sake of reproducibility and comparability, CORD-19, M17+ and Hate Speech sets were split into train and test sets in an 80:20 ratio with a fixed random seed before the application of feature subsetting, while BANKING77 and Food Hazard were already split from the dataset source. This resulted in the same respective instances being tested each time, only with different features.
For all datasets, we experimented with the following feature set versions:
  • BoW only is a feature set where the only features used are bag-of-words terms weighted by their TF-IDF score along with the target variable.
  • LLM-features only dataset has all columns beside the target variable LLM-generated.
  • BoW + LLM-generated features is a dataset with TF-IDF-weighted terms and all LLM-features appended.
Additionally, for two datasets from the scientific domain (CORD-19, M17+), we used the following three additional feature sets for experiments focused on user-specified features:
  • text only datasets we used only title and abstract as input in the AutoGluon Tabular model.
  • SciBERT embeddings only is a feature set where each abstract was embedded using the SciBERT model [4]. SciBERT was trained on a large amount of scientific data and significantly outperformed the base BERT model [10]. The embeddings produced by this model are vectors of length 768.
@@
Detailed settings of all baseline methods can be found on the GitHub page referenced from the Availability of code and data section.
@@ 2.7 Action rules
An action rule $r_{a}$ can be decomposed into two classification rules, $r_{undesired}$ and $r_{desired}$: one that predicts the undesired state of the target (before intervention), and another that predicts the desired state (after intervention), represented as $r_{undesired}\rightarrow r_{desired}$. Uplift is a measure used to predict the incremental response to an action [25], see Equation 7. Unlike commonly used measures such as confidence and lift, the uplift measure reflects the change over the entire dataset. Therefore, even rules with a relatively small value of uplift can be of practical significance.
@@
During the generation of action rules, it is necessary to address the issue of generating a large number of highly similar rules. The dominant action rule approach to reducing action rules is inspired by closed association rules [23]. This concept is analogous to dominant action rules in that closed rules cannot be extended without a loss of support, which can be likened to the reduction in uplift seen in dominant rules. A dominant action rule is one that cannot be further expanded with additional conditions to increase its uplift. For example, the rule $r_{4}$ (see Equation 8) is dominant over the longer rule $r_{5}$. In this specific case, the conditions in rule $r_{4}$ are not extended in $r_{5}$ in a way that increases uplift. Instead, the addition of another condition in $r_{5}$ results in a decreased uplift, demonstrating the dominance of $r_{4}$ over $r_{5}$.